# Approximately-symmetric neural networks for quantum spin liquids

### v1

@@ Abstract
We propose and analyze a family of approximately-symmetric neural networks for quantum spin liquid problems. These tailored architectures are parameter-efficient, scalable, and significantly outperform existing symmetry-unaware neural network architectures. Utilizing the mixed-field toric code model, we demonstrate that our approach is competitive with the state-of-the-art tensor network and quantum Monte Carlo methods. Moreover, at the largest system sizes ($N=480$), our method allows us to explore Hamiltonians with sign problems beyond the reach of both quantum Monte Carlo and finite-size matrix-product states. The network comprises an exactly symmetric block following a non-symmetric block, which we argue learns a transformation of the ground state analogous to quasiadiabatic continuation. Our work paves the way toward investigating quantum spin liquid problems within interpretable neural network architectures.
@@
Figure 1: (a) The approximately-symmetric NQS architecture for a mixed-field toric code model. The network computes the ground-state amplitude, $\psi_{s}=\Omega(\sigma(\chi(s)))$, given an input bit string $s$. The circles (squares) represent edge (plaquette) variables. The convolutional neural networks (CNN) $\chi$ and $\Omega$ consist of between 1-16 layers and 2-16 channels (only 1 layer and 4 channels shown for each) and use normalized $\mathbb{C}$-sigmoid and $\mathbb{C}$-ELU non-linearities, respectively [13]. The non-linear map $\sigma$ imposes invariance on all following layers. For training, $\chi$ is initialized to the identity, while $\Omega$ is randomly initialized. The non-symmetric layers “unfatten” the loop symmetry operators of the quantum spin liquid (purple box) in the spirit of quasi-adiabatic continuation [14]. These symmetries are then enforced exactly in the following layers. (b) The convergence of the energy density (blue) and the non-local Bricmont-Frölich-Fredenhagen-Marcu string order parameter (red) as a function of training time (step size times iteration number), in a regime of the mixed-field toric code model which suffers from the sign problem. (c) Phase diagram of the toric code (Eq. 1) as a function of magnetic field strength, with $h_{y}=0.2$, imposing a sign problem. The phase-transition locations are extracted from finite-size extrapolation of the string order parameter [13]. The red arrow indicates the approximate-to-exact mapping carried out by the non-symmetric block of the network depicted in panel (a).
@@
A tremendous amount of recent attention has focused on a new approach, which utilizes neural networks as the variational ansatze (Fig. 1[22]. The interest in such neural quantum states (NQS), owes in part, to theoretical guarantees of their expressivity [23], which is strictly greater than that of efficiently-contractible tensor networks [24]. Moreover, from a more pragmatic perspective, NQS have achieved state-of-the-art ground state energies in certain archetypal models [25, 26], such as the two-dimensional transverse-field Ising model [25].
The simulation of more exotic and delicate quantum phases, such as spin liquids, remains challenging for both NQS and more traditional methods [27, 4, 28, 29]. Indeed, for neural quantum states, despite recent progress on the $J_{1}$-$J_{2}$ Heisenberg model [30, 31, 32], the long-range entangled nature of quantum spin liquids leads to inherently complicated optimization landscapes [33]. This causes the training of generic network architectures to become trapped in local minima [Fig. 2(b)].
@@
Our main results are threefold. First, to impose approximate symmetries on neural quantum states, we leverage techniques from the field of approximately group-equivariant networks [40, 41]. We modify these constructions for quantum many-body problems, incorporating physical insights into the structure of the neural network. Next, we demonstrate the accuracy of our approach on a paradigmatic quantum spin liquid model: the $\mathbb{Z}_{2}$ toric code perturbed by a magnetic field. We show that the variational energies obtained: (i) outperform conventional NQS methods; (ii) converge to exact diagonalization results for small system sizes [Fig. 2(b)]; (iii) match state-of-the-art tensor network and quantum Monte Carlo results for larger system sizes [Fig. 2(c)]; and (iv) enable access to large system sizes ($N=480$) even when the Hamiltonian has a significant sign problem, beyond the reach of both QMC and finite-size matrix product state methods [Fig. 1(c)]. Finally, we discuss how the approximate-symmetries framework facilitates NQS interpretability. In particular, we argue that the neural network discovers a representation of the emergent ground-state symmetries of spin liquids in the spirit of the quasi-adiabatic continuation of Hastings and Wen [Fig. 1(a)] [14].
@@
Approximate symmetries in neural networks—In order to exploit the emergent symmetries to improve the NQS ansatz, we must first construct an approximately-invariant neural network. For a system composed of $N$ qubits, each with state-space {|±$\{\left|\pm 1\right\rangle\}$, a many-body quantum state vector $|\psi\rangle$ can be decomposed into a complete basis labeled by $2^{N}$ bit strings $s$: $|\psi\rangle=\sum_{s}\psi_{s}|s\rangle$, where $\psi_{s}\in\mathbb{C}$ is a complex amplitude and ${|s\rangle}$ is the quantum state associated with the bit string (e.g. $|s\rangle=|-1,1,1,-1,\cdots\rangle$).
@@
For approximate symmetries, the strict invariance condition above is relaxed to $\mathbb{E}_{g\in G}\mathbb{E}_{s}|\psi_{gs}-\psi_{s}|<\epsilon$, where the expectation value $\mathbb{E}$ is taken over all group elements and input bit strings [48]. Given a fully invariant neural network with $\epsilon=0$, one can lift the strict constraints by adding an extra non-invariant layer to the network or by using a non-invariant skip-connection [41, 40]. In principle, for sufficiently large invariance breaking, $\epsilon$, a neural network constructed in such a fashion can target any vector in the Hilbert space. In practice, the appropriate value of $\epsilon$ is learnt by the network itself, and is independent of network hyperparameters [13].
@@
Approximately-invariant neural quantum states for the toric code—We propose a family of approximately-symmetric neural networks utilizing the so-called “combo” architecture [41, 13]. While we focus on the mixed field toric code model [Eqn. 1], our approach is applicable to a broad class of quantum spin liquid problems.
@@
Our proposed architecture is schematically depicted in Fig. 2(a) and structured as follows: we first impose the constraints $A_{v}|\psi\rangle=|\psi\rangle$ on the neural network and then weakly break these constraints by transforming the input with a non-invariant layer. More specifically, the neural network is defined by $\psi_{s}=\Omega(\sigma(\chi(s)))$. Here, $s$ is the bit string input, $\chi$ is a non-invariant convolutional layer acting on the qubits and $\sigma$ is a $G_{\mathrm{TC}}$-invariant non-linearity which maps qubits to plaquettes. Finally, $\Omega$ is a further convolutional layer (consisting of square-shaped kernels) acting on the plaquettes themselves followed by a summation and exponentiation to calculate $\psi_{s}$ [Fig. 2(a)]. The non-invariant convolutional layer $\chi$ has a kernel centered at each link of the lattice, and explicitly breaks the $G_{\textrm{TC}}$ symmetry. Meanwhile, the non-linear layer, $\sigma$, is constructed using the $G_{\mathrm{TC}}$-invariant operators, $B_{p}=\prod_{j\in p}Z_{j}$ (since $[B_{p},A_{v}]=0$) and ensures the $G_{\mathrm{TC}}$-invariance of any further layers [36, 49]. We describe a number of other choices for approximately-symmetric architectures, along with their implementation details, in the Suppl. Mat. [13].
@@
Figure 2: Benchmarking the approximately-symmetric NQS applied to the sign-problem-free toric code with $(h_{x},h_{y},h_{z})=(0.2,0.0,0.2)$. (a) The convergence of the energy density as a function of training time (step size times iteration number) for a $4\times 4$ ($N=24$) lattice. Restricted Boltzman machines (solid purple) with lattice symmetries become stuck in local minima, while the approximately-symmetric NQS (solid brown) converges to the exact diagonalization result (dashed black). Different shades correspond to different random initializations and network hyperparameters [13]. (b) At larger system sizes, where exact diagonalization is unavailable, we compare our approximately-symmetric NQS to state-of-the-art DMRG (dashed brown) and QMC (dotted brown) calculations. (Inset): Zoom-in comparing the NQS, DMRG and QMC energy densities for $L=10$. For further analysis under different perturbations and with different network hyperparameters, see supplemental materials [13].
@@
Let us begin benchmarking the accuracy of the approximately-symmetric NQS architecture for the mixed-field toric-code Hamiltonian in the sign-problem-free case (with $h_{x}=h_{z}=0.2$ and $h_{y}=0$). Starting with small $4\times 4$ lattices ($N=24$ spins), we compute the ground state energy density, $E_{\textrm{gs}}$, using (i) our architecture; (ii) a conventional restricted Boltzmann machine (RBM), with lattice symmetries implemented; and (iii) exact diagonalization. The energy as a function of training time (measured in units of stochastic reconfiguration stepsize $\gamma$ times iteration number $n$) is depicted in Fig. 2(b). While the RBM becomes stuck in local minima for all runs (with different initializations and hyperparameters) [22, 50], the approximately-symmetric architecture converges to the exact diagonalization energy with a relative error, $\delta E=|E^{\textrm{NQS}}_{\textrm{gs}}-E^{\textrm{ED}}_{\textrm{gs}}|/E^{\textrm{ED}}_{\textrm{gs}}\sim 10^{-8}$.
@@
We further find that the performance of the approximately-symmetric NQS is not limited to small values of the symmetry-violating $h_{z}$ field. Indeed, it is accurate to a relative error of $\sim 10^{-5}$ even outside of the topological phase at $h_{z}=0.7$ [13]. This suggests that the ansatz is more widely applicable than naively expected and enables us to identify the phases and phase transitions out of the spin-liquid state within one NQS architecture.
@@
Figure 3: Detecting phase transitions in the toric code with a sign-problem at $h_{y}=0.3$ [Fig. 1(c)]. (a,c) Depicts the derivative of the BFFM string order parameter $O_{Z}$ ($O_{X}$) with respect to the magnetic field, calculated for a square loop around the center of the lattice with perimeter $\ell=8$ ($\ell=4$). We use finite-size extrapolation of the position of the peaks in order to estimate the location of the phase transition (dashed gray line, extrapolation in [13]). (b,d) The second Renyi entropy for a square, 4-qubit subsystem in the center of the lattice. Panels (a) and (c) illustrate the horizontal cut (at $h_{x}=0$) through the phase diagram, while panels (b) and (d) illustrate the vertical cut (at $h_{z}=0$) through the phase diagram.
@@
To demonstrate the scalability of the architecture, we perform an extensive set of numerical simulations on lattices up to $16\times 16$ ($N=480$ spins) using three methods: state-of-the-art DMRG (density matrix renormalization group) [51, 52, 53, 15], continuous-time QMC [42, 54, 55], and the approximately-symmetric NQS architecture. As illustrated in Fig. 2(c), the approximately-symmetric NQS yields competitive energies at all system sizes [inset, Fig. 2(c)]. Due to memory constraints, we were unable to obtain converged DMRG results for the $16\times 16$ lattice.
@@
As shown in Fig. 3, both of these expectations are borne out by the NQS simulations (at $h_{y}=0.3$). In particular, fixing $h_{x}=0$ and sweeping $h_{z}$, we observe a pronounced peak in the derivative of the string order parameter, $\mathrm{d}O_{z}/\mathrm{d}h_{z}$ [Fig. 3(a)], as well as a sharp change in the entanglement entropy at the same field strength [Fig. 3(b)]. Analogous results fixing $h_{z}=0$ and sweeping $h_{x}$ are depicted in Figs. 3(c,d). To estimate the location of the thermodynamic critical point [Fig. 1(c)], we perform a power-law extrapolation in $1/L$ of the location of the observed finite-size cross-over [13]. After this finite-size extrapolation, the phase boundaries in [Fig. 1(c)] match those obtained by “infinite system-size” iPEPS and PCUT computations to about $5\%$ (Fig. 2 of [61, 62]).
@@
To test this hypothesis, we investigate the following two-step training scheme [Fig. 4(a]): First, we train the network in the fully symmetric regime at $(h_{x},h_{y},h_{z})=(0.2,0.0,0.0)$, by optimizing only the parameters of the invariant $\Omega$ layer and keeping the non-invariant $\chi$ layer fixed to the identity. Second, we fix the invariant $\Omega$ layer at these optimized parameters and turn on a field ($h_{z}=0.2$) which breaks the exact invariance of the state under the group, $G_{\textrm{TC}}$; then, we find the ground state by only optimizing the parameters of the $\chi$ layer. Despite this restricted training procedure, we still obtain an excellent ground state energy [Fig. 4(b)], strongly suggesting that the network is indeed effectively learning the “fattened” loops [Fig. 1(a), red arrow] [13].
@@
Outlook—Our work opens the door to a number of intriguing directions. First, by generalizing our approach to arbitrary groups [63, 64], it may be possible to apply the approximately-symmetric NQS framework to Dirac spin liquids or chiral spin liquids [65]. Second, on the numerical front, there are multiple avenues for further improvement, including: (i) tuning the depth of the architecture, (ii) exploring alternative representations for the complex amplitudes required for models with sign problems, (iii) increasing parameter efficiency with pooling layers, or (iv) replacing the CNNs with more elaborate backbone architectures (e.g. transformers) [26]. Finally, although our focus here is on the toric code spin liquid ground states, the approximately-symmetric architecture can be applied for any model with approximate symmetries, as well as real-time evolution problems.
@@
Acknowledgements We gratefully thank DinhDuy Vu, Andrea Pizzi, Johannes Feldmeier, Quynh Nguyen, Rui Wang, Wen-Tao Xu, Marc Machaczek, Di Luo, Phil Crowley, Yichen Huang, Marc Finzi, Max Welling, Shivaji Sondhi, Saumya Shivam, Ruben Verresen, Robert Huang, Leo Lo, Joaquin Rodriguez-Nieva, Lode Pollet, Arthur Pesah for interesting discussions. We also thank Curtis McMullen for an inspiring group theory class. We acknowledge support from the NSF via the STAQ II program and the QLCI program (grant no. OMA-2016245), and from the Wellcome Leap under the Q for Bio program. D.K. acknowledges support from a Generation-Q AWS and HQI fellowship. S.M.L. acknowledges support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Agreement no 948141) from ERC Starting Grant SimUcQuam. N.Y.Y. acknowledges support from a Simons Investigator Award.
Supplementary Information:
Approximately-symmetric neural networks for quantum spin liquids
Dominik S. Kufel*1,2, Jack Kemp*1,2, Simon M. Linsel 3,4, Chris R. Laumann5, Norman Y. Yao1,2
1Department of Physics, Harvard University, 17 Oxford St., MA 02138, USA
2Harvard Quantum Initiative, 60 Oxford St., MA 02138, USA
3Faculty of Physics, Arnold Sommerfeld Centre for Theoretical Physics (ASC),
Ludwig-Maximilians-Universität München, Theresienstr. 37, 80333 München, Germany
4Munich Center for Quantum Science and Technology (MCQST), Schellingstr. 4, 80799 München, Germany
5Department of Physics, Boston University, 590 Commonwealth Avenue, Boston, Massachusetts 02215, USA
@@
I Neural quantum states and symmetries of many-body systems

### v2

@@ Abstract
We propose and analyze a family of approximately-symmetric neural networks for quantum spin liquid problems. These tailored architectures are parameter-efficient, scalable, and significantly outperform existing symmetry-unaware neural network architectures. Utilizing the mixed-field toric code and PXP Rydberg Hamiltonian models, we demonstrate that our approach is competitive with the state-of-the-art tensor network and quantum Monte Carlo methods. Moreover, at the largest system sizes ($N=480$ for toric code, $N=1584$ for Rydberg PXP), our method allows us to explore Hamiltonians with sign problems beyond the reach of both quantum Monte Carlo and finite-size matrix-product states. The network comprises an exactly symmetric block following a non-symmetric block, which we argue learns a transformation of the ground state analogous to quasiadiabatic continuation. Our work paves the way toward investigating quantum spin liquid problems within interpretable neural network architectures.
@@
Figure 1: (a) The approximately-symmetric NQS architecture for a mixed-field toric code model. The network computes the ground-state amplitude, $\psi_{s}=\Omega(\sigma(\chi(s)))$, given an input bit string $s$. The circles (squares) represent edge (plaquette) variables. The convolutional neural networks (CNN) $\chi$ and $\Omega$ consist of between 1-16 layers and 2-16 channels (only 1 layer and 4 channels shown for each) and use normalized $\mathbb{C}$-sigmoid and $\mathbb{C}$-ELU non-linearities, respectively (Supp. Mat.). The non-linear map $\sigma$ imposes invariance on all following layers. For training, $\chi$ is initialized to the identity, while $\Omega$ is randomly initialized. The non-symmetric layers “unfatten” the loop symmetry operators of the quantum spin liquid (purple box) in the spirit of quasi-adiabatic continuation [13]. These symmetries are then enforced exactly in the following layers. (b) The convergence of the energy density (blue) and the non-local Bricmont-Frölich-Fredenhagen-Marcu string order parameter (red) as a function of training time (step size times iteration number), in a regime of the mixed-field toric code model which suffers from the sign problem. (c) Phase diagram of the toric code (Eq. (1)) as a function of magnetic field strength, with $h_{y}=0.2$, imposing a sign problem. The phase-transition locations are extracted from finite-size extrapolation of the string order parameter (Supp. Mat.). The red arrow indicates the approximate-to-exact mapping carried out by the non-symmetric block of the network depicted in panel (a).
@@
A tremendous amount of recent attention has focused on a new approach, which utilizes neural networks as the variational ansatze (Fig. 1[21]. The interest in such neural quantum states (NQS), owes in part, to theoretical guarantees of their expressivity [22], which is strictly greater than that of efficiently-contractible tensor networks [23]. Moreover, from a more pragmatic perspective, NQS have achieved state-of-the-art ground state energies in certain archetypal models [24, 25], such as the two-dimensional transverse-field Ising model.
The simulation of more exotic and delicate quantum phases, such as spin liquids, remains challenging for both NQS and more traditional methods [26, 4, 27, 28]. Indeed, for neural quantum states, despite recent progress on the $J_{1}$-$J_{2}$ Heisenberg model [29, 30, 31, 32, 33], the long-range entangled nature of quantum spin liquids leads to inherently complicated optimization landscapes [34, 35]. This causes the training of generic network architectures to become trapped in local minima [Fig. 2(a)].
@@
Our main results are threefold. First, to impose approximate symmetries on neural quantum states, we leverage techniques from the field of approximately group-equivariant networks [42, 43]. We modify these constructions for quantum many-body problems, incorporating physical insights into the structure of the neural network. Next, we demonstrate the accuracy of our approach on a paradigmatic quantum spin liquid model: the $\mathbb{Z}_{2}$ toric code perturbed by a magnetic field. We show that the variational energies obtained: (i) outperform conventional NQS methods; (ii) converge to exact diagonalization results for small system sizes [Fig. 2(a)]; (iii) match state-of-the-art tensor network and quantum Monte Carlo results for larger system sizes [Fig. 2(b)]; and (iv) enable access to large system sizes ($N=480$) even when the Hamiltonian has a significant sign problem, beyond the reach of both QMC and finite-size matrix product state methods [Fig. 1(b)]. Finally, we discuss how the approximate-symmetries framework facilitates NQS interpretability. In particular, we argue that the neural network discovers a representation of the emergent ground-state symmetries of spin liquids in the spirit of the quasi-adiabatic continuation of Hastings and Wen [Fig. 1(a)] [13].
@@
Approximate symmetries in neural networks—In order to exploit the emergent symmetries to improve the NQS ansatz, we must first construct an approximately-invariant neural network. For a system composed of $N$ qubits, each with state-space $\{\left|\pm 1\right\rangle\}$, a many-body quantum state vector $|\psi\rangle$ can be decomposed into a complete basis labeled by $2^{N}$ bit strings $s$: $|\psi\rangle=\sum_{s}\psi_{s}|s\rangle$, where $\psi_{s}\in\mathbb{C}$ is a complex amplitude and $|s\rangle$} is the quantum state associated with the bit string (e.g. $|s\rangle=|-1,1,1,-1,\cdots\rangle$).
@@
For approximate symmetries, the strict invariance condition above is relaxed to $\mathbb{E}_{g\in G}\mathbb{E}_{s}|\psi_{gs}-\psi_{s}|<\epsilon$, where the expectation value $\mathbb{E}$ is taken over all group elements and input bit strings [50]. Given a fully invariant neural network with $\epsilon=0$, one can lift the strict constraints by adding an extra non-invariant layer to the network or by using a non-invariant skip-connection [43, 42]. In principle, for sufficiently large breaking $\epsilon$ of the fixed point symmetries, a neural network constructed in such a fashion can target any vector in the Hilbert space. In practice, the appropriate value of $\epsilon$ is learnt by the network itself, and is independent of network hyperparameters (Supp. Mat.).
@@
Approximately-invariant neural quantum states for the toric code—We propose a family of approximately-symmetric neural networks utilizing the so-called “combo” architecture [43]. While we focus on the mixed field toric code model [Eq. (1)], our approach is applicable to a broad class of quantum spin liquid problems.
@@
Our proposed architecture is schematically depicted in Fig. 1(a) and structured as follows: we first impose the constraints $A_{v}|\psi\rangle=|\psi\rangle$ on the neural network and then weakly break these constraints by transforming the input with a non-invariant layer. More specifically, the neural network is defined by $\psi_{s}=\Omega(\sigma(\chi(s)))$. Here, $s$ is the bit string input, $\chi$ is a non-invariant convolutional layer acting on the qubits and $\sigma$ is a $G_{\mathrm{TC}}$-invariant non-linearity which maps qubits to plaquettes. Finally, $\Omega$ is a further convolutional layer (consisting of square-shaped kernels) acting on the plaquettes themselves followed by a summation and exponentiation to calculate $\psi_{s}$ [Fig. 1(a)]. The non-invariant convolutional layer $\chi$ has a kernel centered at each link of the lattice, and explicitly breaks the $G_{\textrm{TC}}$ symmetry. Meanwhile, the non-linear layer, $\sigma$, is constructed using the $G_{\mathrm{TC}}$-invariant operators, $B_{p}=\prod_{j\in p}Z_{j}$ (since $B_{p},A_{v}]=0$) and ensures the $G_{\mathrm{TC}}$-invariance of any further layers [38, 51]. We describe a number of other choices for approximately-symmetric architectures, along with their implementation details, in the  (Supp. Mat.).
@@
Figure 2: Benchmarking the approximately-symmetric NQS applied to the sign-problem-free toric code with $(h_{x},h_{y},h_{z})=(0.2,0.0,0.2)$. (a) The convergence of the energy density as a function of training time (step size times iteration number) for a $4\times 4$ ($N=24$) lattice. Convolutional neural networks (CNNs) (solid purple) become stuck in local minima, while the approximately-symmetric NQS (solid brown) converges to the exact diagonalization result (dashed black). Different shades correspond to different random initializations and network hyperparameters (Supp. Mat.). (b) At larger system sizes, where exact diagonalization is unavailable, we compare our approximately-symmetric NQS to state-of-the-art DMRG (dashed brown) and QMC (dotted brown) calculations. Due to memory constraints, we were unable to obtain converged DMRG results for the $16\times 16$ lattice. (Inset): Zoom-in comparing the NQS, DMRG and QMC energy densities for $L=10$. For further analysis under different perturbations and with different network hyperparameters, see (Supp. Mat.). NQS error bars are shown as shading. QMC uncertainty (not shown) is of order $\sim 10^{-4}$ in units of energy density.
@@
Let us begin benchmarking the accuracy of the approximately-symmetric NQS architecture for the mixed-field toric-code Hamiltonian in the sign-problem-free case (with $h_{x}=h_{z}=0.2$ and $h_{y}=0$). Starting with small $4\times 4$ lattices ($N=24$ spins), we compute the ground state energy density, $E_{\textrm{gs}}$, using (i) our architecture; (ii) a conventional convolutional neural network (CNN) and (iii) exact diagonalization. The energy as a function of training time (measured in units of stochastic reconfiguration stepsize $\gamma$ times iteration number $n$) is depicted in Fig. 2(a). While the CNN becomes stuck in local minima for all runs (with different initializations and hyperparameters) [21, 52], the approximately-symmetric architecture converges to the exact diagonalization energy up to a relative error, $\delta E=|E^{\textrm{NQS}}_{\textrm{gs}}-E^{\textrm{ED}}_{\textrm{gs}}|/E^{\textrm{ED}}_{\textrm{gs}}\sim 10^{-8}$ (see also Fig. S2 in the supplemental material (Supp. Mat.)for further comparisons with full-lattice-symmetrized restricted Boltzmann machines and additional validation of the results).
@@
We further find that the performance of the approximately-symmetric NQS is not limited to small values of the symmetry-violating $h_{z}$ field. Indeed, it is accurate to a relative error of $\sim 10^{-5}$ even outside of the topological phase at $h_{z}=0.7$ (Supp. Mat.). This suggests that the ansatz is more widely applicable than naively expected and, due to the inclusion of a non-invariant block in the architecture, enables us to identify the phases and phase transitions out of the spin-liquid state using a single approximately-symmetric NQS architecture.
@@
Figure 3: Detecting phase transitions in the toric code with a sign-problem at $h_{y}=0.3$ [Fig. 1(c)]. (a,c) Depicts the derivative of the BFFM string order parameter $O_{Z}$ ($O_{X}$) with respect to the magnetic field, calculated for a square loop around the center of the lattice with perimeter $\ell=8$ ($\ell=4$). We use finite-size extrapolation of the position of the peaks in order to estimate the location of the phase transition (dashed gray line, extrapolation in (Supp. Mat.)). (b,d) The second Renyi entropy for a square, 4-qubit subsystem in the center of the lattice. Panels (a) and (c) illustrate the horizontal cut (at $h_{x}=0$) through the phase diagram, while panels (b) and (d) illustrate the vertical cut (at $h_{z}=0$) through the phase diagram.
@@
To demonstrate the scalability of the architecture, we perform an extensive set of numerical simulations on lattices up to $16\times 16$ ($N=480$ spins) using three methods: state-of-the-art DMRG (density matrix renormalization group) [53, 54, 55, 14], continuous-time QMC [44, 56, 57], and the approximately-symmetric NQS architecture. As illustrated in Fig. 2(b), the approximately-symmetric NQS yields competitive energies at all system sizes [inset, Fig. 2(b)].
Crucially, this scalability extends to spin liquid models beyond the mixed-field toric code. As an example setting, which has received widespread recent attention, we utilize our approximately symmetric NQS to explore the so-called PXP model (on the ruby lattice), which describes the physics associated with Rydberg atom arrays in the blockade regime [4, 5, 58, 59, 60]. Similar to the toric code setting, we find that our approximately-symmetric neural network significantly outperforms non-gauge-symmetry aware NQS and yields competitive ground state energies compared to finite DMRG methods (see Fig. 5 in End Matter). Moreover, we are able to reach significantly larger system sizes (up to $N=1584$ Rydberg atoms), which will be crucial for benchmarking the next generation of Rydberg atom array experiments [61]. Further application settings natural for our approximately symmetric NQS are discussed in Table I in the Supplementary Material.
Furthermore, to demonstrate the general utility of our architecture, we consider our approach is not limited to the toric code model but is applicable to a broad class of quantum spin liquid problems.
@@
As shown in Fig. 3, both of these expectations are borne out by the NQS simulations (at $h_{y}=0.3$). In particular, fixing $h_{x}=0$ and sweeping $h_{z}$, we observe a pronounced peak in the derivative of the string order parameter, $\mathrm{d}O_{z}/\mathrm{d}h_{z}$ [Fig. 3(a)], as well as a sharp change in the entanglement entropy at the same field strength [Fig. 3(b)]. Analogous results fixing $h_{z}=0$ and sweeping $h_{x}$ are depicted in Figs. 3(c,d). To estimate the location of the thermodynamic critical point [Fig. 1(c)], we perform a power-law extrapolation in $1/L$ of the location of the observed finite-size cross-over (Supp. Mat.). After this finite-size extrapolation, the phase boundaries in [Fig. 1(c)] match those obtained by “infinite system-size” infinite projected-entangled pair states (iPEPS) and perturbative continuous unitary transformation (pCUT) computations to about $5\%$ (Fig. 2 of [67, 68]).
@@
To test this hypothesis, we investigate the following two-step training scheme [Fig. 4(a]): First, we train the network in the fully symmetric regime at $(h_{x},h_{y},h_{z})=(0.2,0.0,0.0)$, by optimizing only the parameters of the invariant $\Omega$ layer and keeping the non-invariant $\chi$ layer fixed to the identity. Second, we fix the invariant $\Omega$ layer at these optimized parameters and turn on a field ($h_{z}=0.2$) which breaks the exact invariance of the state under the group, $G_{\textrm{TC}}$; then, we find the ground state by only optimizing the parameters of the $\chi$ layer. Despite this restricted training procedure, we still obtain an excellent ground state energy [Fig. 4(b)], strongly suggesting that the network is indeed effectively learning the “fattened” loops [Fig. 1(a), red arrow] (Supp. Mat.).
@@
Outlook— Our work opens the door to a number of intriguing directions. First, straightforward extensions of the approximately-symmetric architecture may be used to study increasingly exotic abelian spin liquid models, ranging from experimentally-motivated microscopic models of $\mathds{Z}_{2}$ liquids [5], to more general chiral and non-chiral $\mathds{Z}_{p}$ and $U(1)$ liquids. By further adapting techniques from [69, 70], models with non-abelian anyons and/or non-abelian continuous symmetries may also be accessible (see (Supp. Mat.)for more discussion).
@@
Second, on the numerical front, there are multiple avenues for further improvement, including: (i) tuning the depth of the architecture, (ii) exploring alternative representations for the complex amplitudes required for models with sign problems, (iii) increasing parameter efficiency with pooling layers, or (iv) replacing the CNNs with more elaborate backbone architectures (e.g. transformers) [25]. Finally, although our focus here has been on finding ground states, the approximately-symmetric architecture can also be applied to simulate real-time dynamics in such systems.
@@
Acknowledgements We gratefully thank Andrea Pizzi, Johannes Feldmeier, Lode Pollet, Quynh Nguyen, Rui Wang, Wen-Tao Xu, Marc Machaczek, Di Luo, Phil Crowley, Yichen Huang, Marc Finzi, Max Welling, Shivaji Sondhi, Saumya Shivam, Ruben Verresen, Robert Huang, Leo Lo, Joaquin Rodriguez-Nieva, Arthur Pesah for interesting discussions. We also thank Curtis McMullen for an inspiring group theory class. We acknowledge support from the NSF via the STAQ program and the QLCI program (grant no. OMA-2016245), and from the Wellcome Leap under the Q4Bio program. D.K. acknowledges support from a Generation-Q AWS and HQI fellowship. S.M.L. acknowledges support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Agreement no 948141) from ERC Starting Grant SimUcQuam. N.Y.Y. acknowledges support from a Simons Investigator Award.
The source code supporting this work is publicly available at https://github.com/dom-kufel/Approximate-Symmetries-TC.
End Matter
Spin liquid of the PXP model on the ruby lattice—We present details associated with our study of the PXP model using approximately symmetric NQS. We begin by introducing the setting. Consider a system of $N$ Rydberg atoms placed on an $L\times(2L+1)$ ruby lattice [$N=3L(2L+1)$] with open boundary conditions (Supp. Mat.)and described by the following Hamiltonian
$H=-\frac{\Omega}{2}\sum_{i}(b_{i}+b_{i}^{\dagger})+\sum_{ij}V_{ij}n_{i}n_{j}-\delta\sum_{i}n_{i}$ (3)
where on-site Hilbert space is $\{|g_{i}\rangle,|e_{i}\rangle\}$, $n_{i}=|e_{i}\rangle\langle e_{i}|$ measures the $i$-th site atom occupation number, and $|e_{i}\rangle=b_{i}^{\dagger}|g_{i}\rangle$, $|g_{i}\rangle=b_{i}|e_{i}\rangle$. In the so-called the PXP limit of this Hamiltonian, the interaction strength is assumed to be infinite for atoms within the blockade radius $R_{b}$ (i.e. $r_{ij}\leq R_{b}=2a$ with $a$ being the shortest distance between two atoms), and zero elsewhere; this strictly prohibits the presence of two excited atoms within a Rydberg blockade radius.
For this PXP model, iDMRG numerics on cylinders have shown that the ground state phase diagram consists of three phases [4]: a trivial phase connected to $\left|g\right\rangle^{\otimes N}$ for $\delta/\Omega\lesssim 1.4$, a $\mathbb{Z}_{2}$ quantum spin liquid phase (QSL) for $1.4\lesssim\delta/\Omega\lesssim 2.0$, and a symmetry-broken valence-bond solid (VBS) phase for $\delta/\Omega\gtrsim 2.0$. The QSL and VBS phases arise from a local constraint permitting exactly one excited atom per lattice vertex when $\delta/\Omega\gtrsim 1/4$. There are exponentially many configurations that fulfill these local (Gauss’ law) constraints. In the QSL phase, the ground state is approximately an equal superposition of all fully packed configurations, while in the VBS phase, the state spontaneously collapses into one such particular configuration.
To explore the PXP model using approximately symmetric NQS, we tailor the Wilson loop operators (which act around each hexagon of the ruby lattice) such that they map one Gauss’-law-satisfying configuration to another [4]. Since this symmetry is only approximate for generic values of $\delta$ (even within the PXP limit), our architecture includes a gauge-invariant block alongside a non-symmetric block. Additional details of the implementation are provided in the supplemental material (Supp. Mat.). We compare the performance of this architecture to both RBMs and finite-cluster DMRG. In all cases, we focus on $\delta=1.6$, which corresponds to a point within the QSL phase.
As depicted in Fig. 5(a), we find that, much as in the mixed-field toric code case, our approximately symmetric NQS yields a nearly three order of magnitude improvement in terms of relative energies (dashed line obtained from exact diagonalization) as compared to an RBM architecture. In Fig. 5(b), we present a comparison of the ground-state energy densities obtained via our approximately symmetric NQS with finite DMRG results. For those system sizes accessible to DMRG ($L=6,8,10$), we observe excellent agreement between NQS and DMRG, with a relative energy difference within $\epsilon_{rel}=10^{-3}$$10^{-4}$. In the same figure, we also show the NQS numerics for $L=16$ (equivalent to $N=1584$ Rydberg atoms) which goes beyond the system sizes that have been investigated using either finite DMRG or QMC [60]. Finally, we note that our approach for studying the PXP limit of the Rydberg Hamiltonian, can also be directly applied to the full long-range $1/r^{6}$ interacting model [5], where the existence of a spin liquid ground state remains under debate [4, 60].
Figure 5: (a) Convergence of the energy density $E_{\mathrm{GS}}=\langle H\rangle/N$ for approximately symmetric neural networks (red curves), compared to symmetrized RBMs (blue curves) and exact diagonalization (dashed black) on a system of $N=30$ Rydberg atoms. Inset: schematic depiction of Rydberg atoms trapped in optical tweezers on a ruby lattice [5]. (b) Energy density convergence for larger system sizes $N=234,408,630,1584$ using approximately symmetric neural quantum states (solid curves), compared to finite-cluster DMRG (dashed black curves) with bond dimension $\chi=1024$ and $100$ sweeps.
Supplementary Material
@@
Appendix S1 Neural quantum states and symmetries of many-body systems